OpenStudy (anonymous):

Write an equation of the line containing the given point and perpendicular to the given line. (8,-3); 8x + 5y=4 The equation of the line is y = ?

OpenStudy (anonymous):

So the original equation in y=mx+b form is: \[y=(-8/5)x + 4\] The slope of the perpendicular line will be the inverse of the negative of this slope. So we can substitute the given point into the following equation: \[y=(5/8)x+b\] Giving us: \[-3=(5/8)8+b\] \[-3=5+b\] b=-8 With a final answer of: \[y=(5/8)x-8\]

OpenStudy (anonymous):

Thank you!